OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 5 Any Real Number No Solution 0 3

OpenStudy (anonymous):

no one cares

mathslover (mathslover):

@ASVPJOHNNY , this is seriously an annoying comment. Please behave nicely here on OpenStudy.

OpenStudy (anonymous):

1D DIZ DIK

OpenStudy (anonymous):

WHO IZ U

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

MATHSLOVER WHO R U

OpenStudy (anonymous):

thanks bless

OpenStudy (anonymous):

@@ ...

mathslover (mathslover):

@ASVPJOHNNY , I am an OpenStudy Ambassador and before that, I am an OpenStudy user.

OpenStudy (anonymous):

@ASVPJOHNNY pls respect

OpenStudy (anonymous):

1D SHE DONT LIKE U

OpenStudy (anonymous):

@OnlyIn1D no problem :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

2(x+1)=2x+5 =>2x+2=2x+5 =>2x-2x=5-2 no solution :)